OpenStudy (anonymous):

select the equation that is perpendicular to y=1/4x+11 and passes through point (3,5)

OpenStudy (texaschic101):

can you tell me the slope of the equation given ?

OpenStudy (anonymous):

sadly no, but and it has to be in slope intercept form

OpenStudy (texaschic101):

in y = mx + b form, the m is the slope. so in y = 1/4x + 11...what is the slope ?

OpenStudy (anonymous):

i was not given the slope. the exact question is stated in the top.

OpenStudy (texaschic101):

you have to find the slope in order to do these types of problems. y = 1/4x + 11.....the 1/4 is in the m position, and is therefore the slope. Now if you needed a parallel line, the slope would stay the same. But you need a perpendicular line, therefore, the slope is the negative reciprocal of the slope in the equation. All that means is " flip " the slope and change the sign. The negative reciprocal of 1/4 = -4/1 or just -4 (see how I flipped the slope and changed the sign) Now we will use y = mx + b and solve for b slope(m) = -4 (3,5) x = 3 and y = 5 now we sub y = mx + b 5 = -4(3) + b 5 = -12 + b 5 + 12 = b 17 = b so your perpendicular equation is : y = -4x + 17 does that make sense to you ? Any questions ?

OpenStudy (anonymous):

oooooooooooh k thanks this really helped a lot

OpenStudy (texaschic101):

glad to help :)